. No Money Bank has an issue of preferred stock with a $6.5 stated dividend that just sold for $90 per share. What is the bank’s cost of preferred stock? What if the floatation cost is 2% of the current market price? (Answer: 7.22%, 7.4%)
  

steps required
Answer #1

PRICE = 90, DIVIDEND = 6.5

COST OF PREFERRED STOCK = DIVIDEND/PRICE = 6.5/90 = 0.07222 = 7.22%

ANS 1 : 7.22%

NOW IF FLOTATION COST = 2% OF MARKET PRICE

f = 90 X 2% = 1.8

COST OF PREFERRED STOCK = DIVIDEND/(PRICE - f)  

COST OF PREFERRED STOCK = 6.5/(90- 1.8) = 6.5/ 88.2 = 0.073696 = 7.4% (rounded to 1 decimal)

ANS 2 : 7.4% (Thumbs up please)